Apply the method of Varition of parameter

d^2y/dx^2 -4 dy/dx +3 y = 1/1+e^-x


1
Expert's answer
2021-10-18T17:10:33-0400

"\frac{d^2y}{dx^2}-4\frac{dy}{dx}+3y=\frac{1}{1+e^{-x}}\\
y''-4y'+3=\frac{1}{1+e^{-x}}\\
\text{The auxilliary equation is; }\\
m^2-4m+3=0\\
(m-1)(m-3)=0\\
m=1 \text{ or } 3\\
\text{The complimentary solution is; }\\
y=Ae^x+Be^{3x}\\
\text{The Wronskian for the DE is;}\\
W=y_1y_2'-y_1'y_2\\
W=3e^{4x}-e^4x=2e^4x\\
C_1'=\frac{-f(x)y_2}{W}=\frac{e^{3x}}{2e^{4x}(1+e^{-x})}=\frac{1}{2e^x(1+e^{-x})}=\frac{1}{2e^x+2}\\
C_2'=\frac{f(x)y_1}{W}=\frac{e^x}{2e^{4x}(1+e^{-x})}=\frac{1}{(2e^{3x}+2e^{2x})}\\
C_1=\int \frac{1}{2e^x(1+e^{-x})} dx=\frac{1}{2}(x-\ln(e^x+1))+K_1\\
C_2=\int \frac{1}{(2e^{3x}+2e^{2x})} dx= \frac{1}{4}(2x-e^{-2x}+2e^{-x}-2\ln(e^x+1))+K_2\\
\text{Therefore, }\\
y(x)=Ae^x+Be^{3x}+\frac{e^x}{2}(x-\ln(e^x+1))+\frac{e^{3x}}{4}(2x-e^{-2x}+2e^{-x}-2\ln(e^x+1))+K"